61) Which of these assets were overvalued (in a bubble) before Great Recession hit the US economy?

Crude Oil
US Treasury bonds
Real Estate
Silver

62)

If investment falls by $5 billion. The marginal propensity to consume is 0.75. So Real GDP

Increases by $20 billion
Decreases by $20 billion
Increases by $40 billion

Decreases by $25 billion

65)

If the marginal propensity to consume is 0.8 and there is a $20 million increase in investment spending, then the aggregate demand curve will shift

to the left by $7.5 million
to the right by:$140 million
to the left by $40 million.
to the right by:$100 million

67)

The $787 billion "stimulus plan" passed by Congress in 2009 is an example of

monetary policy
tax rate policy
interest rate policy
fiscal policy
Answer #1

a) "C"

the real state witnessed the bubble before the great recession.

b) If the MPC is 0.75 the multiplier is 1/1-0.75 = 1/ 0.25 = 4. The GDP will decrease by 5 x 4 = 20 billion. the answer is "B" decrease by 20 billion.

c) If the MPC is 0.8 the multiplier will be 5. an increase in the investment will result in total increase of 20x 5 = 100 million. the answer is "D".

d) It Is an example of "D" expansionary fiscal policy.